multiple choice physics question

What is the magnitude and direction of the electric force on a -1.2 µC charge at a point where the electric field is 2500 N/C and is directed along the +y axis.

0.15 N, +y direction
0.15 N, -y direction
4.3 N, +x direction
0.0030 N, -y direction
Answer #1
Given:charge q = -1.2 *10^-6 Celectric field E = 2500 N/Celectric force F = Eq= 2500*(-1.2*10^-6)=0.0030 Ndirection is -y directionso choice is 4.
answered by: ammu